VIBRATIONAL PREDISSOCIATION OF THE HF-HCl COMPLEXES AT ν_{HF} = 3 LEVEL

\underline{ZHENHONG YU}, PATRICK MEDLEY, BRENDAN CONNORS and WILLIAM KLEMPERER, Harvard University, Department of Chemistry and Chemical Biology, Cambridge, Massachusetts 02138.

The binary complex between HF and HCl at ν_{HF} = 3 level, which has two stable forms: HF-HCl and HCl-HF, is studied recently by molecular beam laser induced fluorescence with an intracavity Ti-sapphire ring laser^{a} and the detection of HF Δν = 2 ← 0 emission. Similar to previously reported K = 0-0 (300000) ← (000000) transition of HF-HCl^{b}, the K = 1-0 subband seen at 11353.2 cm^{−1} shows significant rotational dependence to the predissociation linewidths, varying from 1.3 to 2.4 GHz. The rotational A constant of 43.5 cm^{−1} clearly indicates that the vibration-averaged angle between the HF axis and a-axis decreases over valence excitation, compared to 40.1 cm^{−1} at ν_{HF} = 1 level. Another weaker Q-branch band observed at 11351.7 cm^{−1}, with slightly narrower linewidths and little apparent J dependence, is tentatively assigned to the K = 1-0 (301000) ← (000000) combination transition of HCl-HF. The 2D vibration-averaged ab initio interaction potential of the complex predicts the HCl-HF form becomes more stable than HF-HCl through the valence excitation. The isotopic variation^{c} of predissociation lifetime for the observed transitions is still under investigation.

aH. -C. Chang and W. Klemperer, J. Chem. Phys. 98, 2497(1993).

bR. Mollaaghababa, H. -C. Chang and W. Klemperer, Can. J. Phys. 72, 963(1993).

cG. T. Fraser and A. S. Pine, J. Chem. Phys. 91, 637(1989).